OpenStudy (anonymous):

Find the distance between (0, 1) and (8, 11). A. 15 B. 12.8 C. 3.16 D. 19

jimthompson5910 (jim_thompson5910):

use the distance formula d = sqrt( (x2 - x1)^2 + (y2 - y1)^2 )

jimthompson5910 (jim_thompson5910):

in this case (x1,y1) = (0,1) so x1 = 0 and y1 = 1 and (x2,y2) = (8,11) so x2 = 8, y2 = 11

OpenStudy (anonymous):

:( i no good at this

jimthompson5910 (jim_thompson5910):

that's ok, that's why you practice

jimthompson5910 (jim_thompson5910):

d = sqrt( (x2 - x1)^2 + (y2 - y1)^2 ) d = sqrt( (8 - 0)^2 + (11 - 1)^2 ) d = sqrt( 8^2 + 10^2 ) d = ???

OpenStudy (aonz):

practice makes perfect lol :D